# The Fox-Wolfram Moment of jet production in relativistic heavy ion collisions

### v2

@@ III $\text{FWM}_{\text{1st}}^{\text{T}}$ in Pb+Pb collisions
We investigate $H_{1}^{T}$, which is the lower-order member of the set of FWMs related to QCD geometric patterns, and apply it to the field of jet quenching physics. This application allows us to examine how the medium modifies jet event shape of varying multiplicities ($n_{\text{jet}}$). The results indicate that $n_{\text{jet}}=2$ events undergo a pronounced broadening modification due to the influence of nuclear matter. Meanwhile, the event shape for $n_{\text{jet}}>2$ becomes broader due to medium modification, except in the highest $\ln(H_{1}^{T})$ bin where the jet number effect is significant. Additionally, we investigate the underlying reasons for the medium modification of $H_{1}^{T}$ spectra for total jet events, without distinguishing between different multiplicities. This is achieved by analyzing the relative fraction of $H_{1}^{T}$ between the two types of events.
Acknowledgments: We extend our sincere gratitude to Jin-Wen Kang, Wei Dai, Shi-Yong Chen, Meng-Quan Yang, Sa Wang, and Shan-Liang Zhang for their invaluable discussions. We would also like to express our appreciation to the open-source software packages Sherpa Gleisberg et al. (2009), FastJet Cacciari et al. (2012), HepMC2 Dobbs and Hansen (2001) and ROOT Brun and Rademakers (1997). This work has been supported by the Guangdong Major Project of Basic and Applied Basic Research under Grant No. 2020B030103008, as well as the Natural Science Foundation of China with Project Nos. 11935007 and 12035007.

### v4

@@ III $\text{FWM}_{\text{1st}}^{\text{T}}$ in Pb+Pb collisions
Now we consider the $H_{1}^{T}$ distribution for the combination of $n_{\text{jet}}=2$ and $n_{\text{jet}}>2$ events in proportion, as shown in FIG. 5. The nuclear modification factor exhibits a similar pattern to that of $n_{\text{jet}}=2$ events, characterized by suppression at low $H_{1}^{T}$ values and enhancement at high $H_{1}^{T}$ values. To understand why the nuclear medium modification for total ($n_{\text{jet}}\geq 2$) events is as shown in FIG. 5, we draw FIG. 6, which illustrates the relative contributions of $n_{\text{jet}}=2$ and $n_{\text{jet}}>2$ events to $H_{1}^{T}$ in both p+p and Pb+Pb collisions at $\sqrt{s}=5.02$ TeV. In the low $H_{1}^{T}$ region, $n_{\text{jet}}>2$ events constitute a slightly larger proportion than $n_{\text{jet}}=2$ events. As $H_{1}^{T}$ increases, the percentage of $n_{\text{jet}}=2$ ($n_{\text{jet}}>2$) events gradually rises from approximately 40% (60%) in low the $H_{1}^{T}$ region to around 90% (10%) in the high $H_{1}^{T}$ region in p+p collisions. Similarly, the corresponding percentages transition from approximately 46% (54%) to around 89% (11%) in Pb+Pb collisions. The transition in the relative contributions between $n_{\text{jet}}=2$ and $n_{\text{jet}}>2$ events occurs at around $\ln(H_{1}^{T})=-4$. The nuclear modification remains suppressed in the low $H_{1}^{T}$ region, where both $n_{\text{jet}}=2$ and $n_{\text{jet}}>2$ events have similar fractions. In contrast, an enhancement effect is observed in the high $H_{1}^{T}$ region, consistent with the behavior of $n_{\text{jet}}=2$ events, which dominate in this region.
Although the focus of this work is on the study of jet event shapes described by $H_{1}^{T}$, rather than the transverse momentum or energy asymmetry, the latter nonetheless defines the former. We also discuss the medium modification of the transverse momentum asymmetry for $n_{\text{jet}}=2$ events in Appendix B.
@@
We investigate $H_{1}^{T}$, which is the lower-order member of the set of FWMs related to QCD geometric patterns, and apply it to the field of jet quenching physics. This application allows us to examine how the medium modifies jet event shape of varying multiplicities ($n_{\text{jet}}$). The results indicate that $n_{\text{jet}}=2$ events undergo a pronounced broadening modification due to the influence of nuclear matter. Meanwhile, the event shape for $n_{\text{jet}}>2$ becomes broader due to medium modification, except in the highest $\ln(H_{1}^{T})$ bin where the jet number reduction effect is significant. Additionally, we investigate the underlying reasons for the medium modification of $H_{1}^{T}$ spectra for total jet events, without distinguishing between different multiplicities. This is achieved by analyzing the relative fraction of $H_{1}^{T}$ between the two types of events.
Acknowledgments: We extend our sincere gratitude to Jin-Wen Kang, Wei Dai, Shi-Yong Chen, Meng-Quan Yang, Sa Wang, and Shan-Liang Zhang for their invaluable discussions. We would also like to express our appreciation to the open-source software packages Sherpa [118], FastJet [139], HepMC2 [141] and ROOT [142]. This work has been supported by the Guangdong Major Project of Basic and Applied Basic Research under Grant No. 2020B030103008, as well as the Natural Science Foundation of China with Project Nos. 11935007 and 12035007.